11. Two banks are offering interest rates on savings accounts with the following information. Bank A: rate of 5.25% compounded semi-annually. Bank B: rate of 5.10% compounded monthly. Calculate the annual percentage rate (APR) for each and determine the best investment for the individual.
Answer #1


At APR 5.25%
EAR of Bank A =(1+5.25%/2)^2-1 =5.32%
At APR 5.10%
EAR of Bank B =(1+5.10%/12)^12-1=5.22%

The best investment is of Bank A =5.32%